### v1

# Relational Graph Transformer

Abstract
Relational Deep Learning (RDL) is a promising approach for building state-of-the-art predictive models on multi-table relational data by representing it as a heterogeneous temporal graph. However, commonly used Graph Neural Network models suffer from fundamental limitations in capturing complex structural patterns and long-range dependencies that are inherent in relational data. While Graph Transformers have emerged as powerful alternatives to GNNs on general graphs, applying them to relational entity graphs presents unique challenges: (i) Traditional positional encodings fail to generalize to massive, heterogeneous graphs; (ii) existing architectures cannot model the temporal dynamics and schema constraints of relational data; (iii) existing tokenization schemes lose critical structural information. Here we introduce the Relational Graph Transformer (RelGT), the first graph transformer architecture designed specifically for relational tables. RelGT employs a novel multi-element tokenization strategy that decomposes each node into five components (features, type, hop distance, time, and local structure), enabling efficient encoding of heterogeneity, temporality, and topology without expensive precomputation. Our architecture combines local attention over sampled subgraphs with global attention to learnable centroids, incorporating both local and database-wide representations. Across 21 tasks from the RelBench benchmark, RelGT consistently matches or outperforms GNN baselines by up to 18%, establishing Graph Transformers as a powerful architecture for Relational Deep Learning [^1] https://github.com/snap-stanford/relgt.
1 Introduction
Real-world enterprise data, such as financial transactions, supply chain data, e-commerce records, product catalogs, customer interactions, and electronic health records, are predominantly stored in relational databases [8]. These databases typically consist of multiple tables, each dedicated to different entity types, interconnected through primary and foreign key links. This abstraction underpins large quantities of complex, dynamically updated data that scale with business volume, storing potentially immense, unexploited knowledge [13]. However, extracting predictive patterns from such data has traditionally depended on manual feature engineering within complex machine learning pipelines, requiring the transformation of multi-table records into flat feature vectors suitable for models like deep neural networks and decision trees [6].
Relational Deep Learning. To enable end-to-end deep learning, relational databases can be represented as relational entity graphs [13], where nodes correspond to entities and edges capture primary-foreign key relationships. This graph-based representation allows Graph Neural Networks (GNNs) to learn abstract features directly from the underlying data structure, effectively modeling complex dependencies for various downstream prediction tasks. With this setup, which is termed as Relational Deep Learning (RDL), GNNs reduce or eliminate the need for manual feature engineering and often lead to improved performance [42], at a fraction of the traditional model development cost.
Existing gaps. Despite their effectiveness, standard message-passing GNN architectures [15, 29, 18, 48] have notable limitations, such as insufficient structural expressiveness [52, 37, 34] and restricted long-range modeling capabilities [1]. For example, consider an e-commerce database with three tables: customers, transactions, and products, which can be represented as a relational entity graph as in Figure 1. In a standard GNN, transactions are always two hops away from each other, connected only through shared customers. This creates an information bottleneck: transaction-to-transaction patterns require multiple layers of message passing, while product relationships remain entirely indirect in shallow networks. Furthermore, products would never directly interact in a two-layer GNN [42], as their messages must pass through both a transaction and a customer, highlighting the inherent structural constraints of GNN architectures that restrict capturing long-range dependencies.
Graph Transformers (GTs) have emerged as more expressive models for graph learning, utilizing self-attention in the full graph to increase the range of information flow and additionally, incorporating positional and structural encodings (PEs/SEs) to better capture graph topology [9, 53, 41]. These advances have produced strong results across domains [38], including foundation models for molecular graphs [46]. However, many GT designs are limited to non-temporal, homogeneous, and small-scale graphs, assumptions that do not hold for relational entity graphs (REGs) [13], which are typically (i) heterogeneous, with different tables representing distinct node types; (ii) temporal, with entities often associated with timestamps and requiring careful handling to prevent data leakage; (iii) large-scale, containing millions or more records across multiple interconnected tables. In particular, existing PEs often require precomputation, depend on graph size, and typically do not scale well to large, heterogeneous, or dynamic graphs [3, 26]. For instance, node2vec [17], while more efficient than Laplacian or random walk PEs, can become prohibitively expensive and impractical to compute on massive graphs [40]. These limitations, along with the inability to capture the multi-dimensional complexity of relational structures, render current GTs inadequate for relational databases.
Figure 1: Overview of the RelGT architecture. First, the input relational entity graph (REG) is converted into tokens where each training seed node (such as the customer node in this example) gets a fixed number of neighboring nodes, which are encoded with a multi-element tokenization strategy. These tokens are then passed through a Transformer network that builds both local and global representations, which are then fed to downstream prediction layers.
Present work. We introduce the Relational Graph Transformer (RelGT), the first Graph Transformer specifically designed for relational entity graphs. RelGT addresses key gaps in existing methods by enabling effective graph representation learning within the RDL framework. It is a unified model that explicitly captures the temporality, heterogeneity, and structural complexity inherent to relational graphs. We summarize the architecture as follows (Figure 1):
  • Tokenization: We develop a multi-element tokenization scheme that converts each node into structurally enriched tokens. By sampling fixed-size subgraphs as local context windows and encoding each node’s features, type, hop distance, time, and local structure, RelGT captures fine-grained graph properties without expensive precomputation at the subgraph or graph level.
  • Attention: We develop a transformer network that combines local and global representations, adapting existing GT architectures [41]. The model extracts features from the local tokens while simultaneously attending to learnable global tokens that act as soft centroids, effectively balancing fine-grained structural modeling with database-wide patterns [30].
  • Validation: We showcase RelGT’s effectiveness through a comprehensive evaluation on 21 tasks from RelBench [42]. RelGT consistently outperforms GNN baselines, with gains of up to 18%, establishing transformers as a powerful architecture for relational deep learning. Compared to HGT [20], a strong GT baseline for heterogeneous graphs, RelGT achieves better results without added computational cost, even when HGT uses Laplacian eigenvectors for positional encoding.
2 Background
2.1 Relational Deep Learning
Relational Deep Learning (RDL) is an end-to-end representation learning framework that converts relational databases into graph structures, enabling neural networks to be applied directly and eliminating the need for manual feature extraction in multi-table data pipelines [13].
Definitions. Formally, we can define a relational database as the tuple $(T,R)$ comprising a collection of tables $T=\{T_{1},\ldots,T_{n}\}$ connected through inter-table relationships $R\subseteq T\times T$. A link $(T_{\text{fkey}},T_{\text{pkey}})\in R$ denotes a foreign key in one table referencing a primary key in another. Each table contains entities (rows) $\{v_{1},\ldots,v_{n_{T}}\}$, with each entity typically consisting of: (1) a unique identifier (primary key), (2) references to other entities (foreign keys), (3) entity-specific attributes, and (4) timestamp information indicating when the entity was created or modified. The structure of relational databases inherently forms a graph representation, called as relational entity graphs (REGs). An REG is formally defined as a heterogeneous temporal graph $G=(V,E,\phi,\psi,\tau)$, where nodes $V$ represent entities from the database tables, edges $E$ represent primary-foreign key relationships, $\phi$ maps nodes to their respective types based on source tables, $\psi$ assigns relation types to edges, and $\tau$ captures the temporal dimension through timestamps [13].
Challenges. Relational entity graphs exhibit three distinctive properties that set them apart from conventional graph data. First, their structure is fundamentally schema-defined, with topology shaped by primary-foreign key relationships rather than arbitrary connections, creating specific patterns of information flow that require specialized modeling approaches. Second, they incorporate temporal dynamics, as relational databases track events and interactions over time, necessitating techniques like time-aware neighbor sampling to prevent future information from leaking into past predictions. Third, they display multi-type heterogeneity, as different tables correspond to different entity types with diverse attribute schemas and data modalities, presenting challenges in creating unified representations that effectively integrate information across diverse node and edge types [44, 49]. These characteristics create both challenges and opportunities for GNN architectures, requiring models that can simultaneously address temporal evolution, heterogeneous information, and schema-constrained structures while processing potentially massive multi-table datasets.
2.2 RDL Methods
The baseline GNN approach introduced by [42] for RDL uses a heterogeneous GraphSAGE [18] model with temporal-aware neighbor sampling, which demonstrates significant improvements compared to traditional tabular methods like LightGBM [27] across all tasks in the RelBench benchmark. This baseline architecture leverages PyTorch Frame’s multi-modal feature encoders [19] to transform diverse entity attributes into initial feature embeddings that serve as input to the GNN. Several specialized architectures have been developed to address specific challenges in relational entity graphs. RelGNN [5] introduces composite message-passing with atomic routes to facilitate direct information exchange between neighbors of bridge and hub nodes, commonly found in relational structures. Similarly, ContextGNN [55] employs a hybrid approach, combining pair-wise and two-tower representations, specifically optimized for recommendation tasks in RelBench.
Beyond pure GNN approaches, retrieval-augmented generation techniques [51] and hybrid tabular-GNN methods [32] have also demonstrated comparable or slightly superior performance to the standard GNN baseline, while showing the use of LLMs [16] and inference speedups, respectively. These approaches confirm the effectiveness of graph, tabular, and LLM-based methods for downstream predictions in relational databases. However, these methods typically optimize specific aspects of the problem, failing to incorporate broader advances from GTs in general graph learning.
2.3 Graph Transformers
Graph Transformers extend the self-attention mechanism from sequence modeling [47] to graph-structured data, offering powerful alternatives to traditional GNNs [9]. These models typically restrict attention to local neighborhoods, functioning as message-passing networks with attention-based aggregation [24, 2], while positional encodings are developed based on Laplacian eigenvectors [10]. Subsequent Graph Transformers incorporate global attention mechanisms, allowing all nodes to attend to one another [53, 36, 31]. This moves beyond the local neighborhood limitations of standard GNNs [1], albeit at the cost of significantly increased computational complexity.
Modern GT architectures have improved the aforementioned early works by creating effective structural encodings and ensuring scalability to medium and large-scale graphs. For structural expressiveness of the node tokens, several positional and structural encoding methods have been developed [12, 3, 33, 22, 26] to inject the input graph topology. For scalability, various strategies have emerged including hierarchical clustering that coarsens graphs [57, 59], sparse attention mechanisms that reduce computational cost [41, 45], and neighborhood sampling techniques for processing massive graphs [58, 4, 30, 11]. Models like GraphGPS [41] combine these advances through hybrid local-global designs that maintain Transformers’ global context advantages while ensuring practical efficiency when scaling to medium and large graph datasets. However, these approaches exhibit several key limitations: they are largely confined to static graphs, and lack mechanisms to handle multiple node and edge types. While specialized Transformers for heterogeneous graphs exist [20, 35, 59, 56], integrating them, alongside other aforementioned methods, into the RDL pipeline remains challenging. This is primarily because adapting positional encodings under precomputation constraints is difficult, compounded by the complexity of modeling large-scale, temporal, and heterogeneous relational entity graphs (REGs).
3 RelGT: Relational Graph Transformer
3.1 Tokenization
Traditional Transformers in natural language processing represent text through tokens with two primary elements: (i) token identifiers (or features) that denotes the token from a vocabulary set and (ii) positional encodings that represent sequential structure [47]. For example, a token can correspond to a word and its positional encoding can correspond to its order in the input sentence. Similarly, Graph Transformers generally adapt this two-element representation to graphs, where nodes are tokens with features, and graph positional encodings provide structural information [9, 28, 41]. Although this two-element approach works well for homogeneous static graphs, it becomes computationally inefficient when trying to encode multiple aspects of graph structural information for REGs.
In particular, capturing heterogeneity, temporality, and schema-defined structure (as defined in Section 2.1) through a single positional encoding scheme would either require complex, multi-stage encoding or result in significant information loss about the rich relational context. For instance, if we were to extend existing PEs for REGs, several practical challenges emerge: (i) standard Laplacian or random walk-based PEs would need significant modification to differentiate between multiple node types (e.g., customers vs. products vs. transactions), (ii) these encodings lack mechanisms to incorporate temporal dynamics critical for time-sensitive predictions (e.g., capturing that a user’s recent purchases are more relevant than older ones), and (iii) the scale of relational databases makes global PE computation in REGs prohibitively expensive. With millions of records across tables, precomputation would only be feasible on small subgraphs, resulting in incomplete structural context.
3.1.1 Proposed Approach
RelGT overcomes these limitations through a multi-element token representation approach, without any computational overhead concerning the dependency on the number of nodes in the input REG. Rather than trying to compress all structural information into a single positional encoding, we decompose the token representation into distinct elements that explicitly model different aspects of relational data. This decoupled design allows each component to capture a specific characteristic of REGs: node features represent entity attributes, node types encode table-based heterogeneity, hop distance preserves relative distances among nodes in a local context, time encodings capture temporal dynamics, and GNN-based positional encodings preserve local graph structure.
[…]
Figure 2: The tokenization procedure. A temporal-aware subgraph sampling step extracts a fixed set of local tokens for each training seed node, denoted by the node in black. Each token incorporates its respective graph structure information, which are element-wise transformed to a common embedding space and combined to form the effective token representation to be fed to the Transformer network.
Encoders. Each element in the 5-tuple is processed by a specialized encoder before being combined into the final token representation, as illustrated in Figure 2.
1. Node Feature Encoder. The node features $x_{v_{j}}$, representing the columnar attributes of the node $v_{j}$ in REG (which corresponds to a table row in a database), are encoded into a $d$-dimensional embedding. Each modality, such as numerical, categorical, multi-categorical, text, and image data, is encoded separately using modality-specific encoders following [19], and the resulting representations are then aggregated into a unified $d$-dimensional embedding.
$h_{\text{feat}}(v_{j})=\text{MultiModalEncoder}(x_{v_{j}})\in\mathbb{R}^{d}$ (1)
[…]
2. Node Type Encoder. The node type encoding steps converts each table-specific entity type $\phi(v_{j})$ to a $d$-dimensional representation, incorporating the heterogeneous information from the input data.
$h_{\text{type}}(v_{j})=W_{\text{type}}\cdot\text{onehot}(\phi(v_{j}))\in\mathbb{R}^{d}$ (2)
where $\phi(v_{j})$ is the node type of $v_{j}$, $W_{\text{type}}\in\mathbb{R}^{d\times|T|}$ is the learnable weight matrix, $|T|$ is the number of node types, and $\text{onehot}(\cdot)$ is the one-hot encoding function.
3. Hop Encoder. The relative hop distance $p(v_{i},v_{j})$, that captures the structural proximity between the seed node $v_{i}$ and a neighbor node $v_{j}$, is encoded into a $d$-dimensional embedding as:
$h_{\text{hop}}(v_{i},v_{j})=W_{\text{hop}}\cdot\text{onehot}(p(v_{i},v_{j}))\in\mathbb{R}^{d}$ (3)
with $p(v_{i},v_{j})$ being the relative hop distance between seed node $v_{i}$ and neighbor node $v_{j}$, and $W_{\text{hop}}\in\mathbb{R}^{d\times h_{\text{max}}}$ the learnable matrix mapping hop distances (up to $h_{\text{max}}$).
4. Time Encoder. The time encoder linearly transforms the time difference $\tau(v_{j})-\tau(v_{i})$ between a neighbor node $v_{j}$ and the seed node $v_{i}$:
$h_{\text{time}}(v_{i},v_{j})=W_{\text{time}}\cdot(\tau(v_{j})-\tau(v_{i}))\in\mathbb{R}^{d}$ (4)
where $\tau(v_{j})-\tau(v_{i})$ is the relative time difference, and $W_{\text{time}}\in\mathbb{R}^{d\times 1}$ are learnable parameters.
5. Subgraph PE Encoder. Finally, for capturing local graph structure that can otherwise not be represented by other elements of the token, we apply a light-weight GNN to the subgraph. This GNN encoder effectively preserves important structural relationships, such as complex cycles and quasi-cliques between entities [25], as well as parent-child relationships (e.g., a product node within the local subgraph corresponding to specific transactions), and can be written as:
$h_{\text{pe}}(v_{j})=\text{GNN}(A_{\text{local}},Z_{\text{random}})_{j}\in\mathbb{R}^{d}$ (5)
where $\text{GNN}(\cdot,\cdot)_{j}$ is a light-weight GNN applied to the local subgraph yielding the encoding for node $v_{j}$, $A_{\text{local}}\in\mathbb{R}^{K\times K}$ is the adjacency matrix of the sampled subgraph of $K$ nodes, and $Z_{\text{random}}\in\mathbb{R}^{K\times d_{\text{init}}}$ are randomly initialized node features for the GNN (with $d_{\text{init}}$ as the initial feature dimension).
One key advantage of using random node features in this GNN encoder is that it breaks structural symmetries between the subgraph topology and node attributes, thereby increasing the expressive power of GNN layers [43]. However, a fixed random initialization would destroy permutation equivariance which is a critical property for generalization. To address this challenge, we resample $Z_{\text{random}}$ independently at every training step. This ‘stochastic initialization’ approach can be viewed as a relaxed version of the learnable PE method described in [26], thus approximately preserving permutation equivariance while retaining the expressivity gains afforded by the randomization.
At last, the effective token representation is formed by combining all encoded elements:
${h}_{\text{token}}(v_{j})=O\cdot\left[{h}_{\text{feat}}(v_{j})\,||\,{h}_{\text{type}}(v_{j})\,||\,{h}_{\text{hop}}(v_{i},v_{j})\,||\,{h}_{\text{time}}(v_{i},v_{j})\,||\,{h}_{\text{pe}}(v_{j})\right]$ (6)
where ||||| | denotes the concatenation of the individual encoder outputs, and $O\in\mathbb{R}^{5d\times d}$ is a learnable matrix to mix the embeddings. This multi-element approach provides a comprehensive token representation that explicitly captures node features, type information, structural position, temporal dynamics, and local topology without requiring expensive computation on the graph structure.
Figure 3: The Transformer network which processes the input tokens by first building local representations using the local tokens, then incorporating global context by attending to centroids that are dynamically updated during training. The final node representations combine both local structural details and global database context, enabling effective prediction across downstream tasks.
3.2 Transformer Network
The Transformer network in RelGT, shown in Figure 3, processes the tokenized relational entity graph using a combination of local and global attention mechanisms, following the successful design components used in modern GTs [41, 50, 30, 11].
Local module. The local attention mechanism allows each seed node to attend to its $K$ local tokens selected during tokenization, capturing the fine-grained relationships defined by the database schema. This mechanism is different from a GNN used in RDL [42] in two key aspects: self-attention is used as the message-passing scheme and the attention is all-pair, i.e., all nodes in the local K set attend to each other. This is implemented using an L layer Transformer network [47] and provides a broader structural coverage compared to a baseline GNN [42]. A practical application of this improvement can be seen in the e-commerce example introduced in Section 1, where the proposed full-attention mechanism can directly connect seemingly unrelated products by identifying relationships through shared transactions or customer behaviors. This capability enables the model to capture subtle associations, such as customers frequently purchasing unexpected combinations of items.
The local node representation $h_{\text{local}}(v_{i})$ is obtained as:
$h_{\text{local}}(v_{i})=\text{Pool}(\text{FFN}(\text{Attention}(v_{i},\{v_{j}\}_{j=1}^{K}))_{L})$ (7)
where, $L$ denotes the layers, FFN and Attention are standard components in a Transformer [47], and Pool denotes the aggregation of $\{v_{j}\}_{j=1}^{K}$ and $v_{i}$ using a learnable linear combination.
Global module. The global attention mechanism enables each seed node to attend to a set of $B$ global tokens representing centroids of all nodes in the graph, conceptually and is adapted from prior works [30, 11]. These centroids are updated during training using an Exponential Moving Average (EMA) K-Means algorithm applied to seed node features in each mini-batch, providing a broader contextual view beyond the local neighborhood. The global representation is formulated as:
$h_{\text{global}}(v_{i})=\text{Attention}(v_{i},\{c_{b}\}_{b=1}^{B})$ (8)
The final output representation of each node $v_{i}$ is obtained by combining local and global embeddings:
$h_{\text{output}}(v_{i})=\text{FFN}([h_{\text{local}}(v_{i})\,||\,h_{\text{global}}(v_{i})])$ (9)
with FFN being a feed forward network. The components of the Transformer in all stages follow standard instantiations with normalization and residual connections.
For downstream prediction, the combined representation of the seed node is passed through a task-specific prediction head. The model is trained end-to-end using suitable task specific loss functions. By leveraging multi-element token representations within a hybrid local-global Transformer architecture, RelGT effectively addresses the challenges of heterogeneity, temporal dynamics, and schema-defined structures inherent in relational entity graphs.
4 Experiments
RelGT is evaluated on the recently introduced Relational Deep Learning Benchmark (RelBench) [42]. RelBench consists of 7 datasets from diverse relational database domains, including e-commerce, clinical records, social networks, and sports, among others. These datasets are curated from their respective source domains and consist a wide range of sizes, from 1.3K to 5.4M records in the training set for the prediction tasks, with a total of 47M training records. For each dataset, multiple predictive tasks are defined, such as predicting a user’s engagement with an advertisement within the next four days or determining whether a clinical trial will achieve its primary outcome within the next year. In total, RelBench has 30 tasks across the 7 datasets, covering entity classification, entity regression, and recommendation. For our evaluation, we focus on 21 tasks on entity classification and regression  [^3] We exclude recommendation tasks in this work since they involve specific considerations, such as identifying target nodes [54] or using pair-wise learning architectures [55] and using RelGT trivially in RDL is sub-optimal..
4.1 Setup and Baselines
We implement RelGT within the RDL pipeline [42] by replacing the original GNN component, while preserving the learning mechanisms, database loaders, and task evaluators. The model has between 10-20 million parameters, and we use a learning rate of $41e-$ 4. For tasks with fewer than one million training nodes, we tune the number of layers $L\in{1,4,8}$ and use dropout rates of ${0.3,0.4,0.5}$. For tasks with more than one million training nodes, we fix the number of layers to $L=4$ due to compute budgets. For the sampling during the token preparation stage, we use $K=300$ local neighbors and set $B=4096$ as the number of tokens for global centroids. For smaller datasets (under one million training nodes), we use a batch size of 256 to ensure sufficient training steps. For larger datasets, we use a batch size of 1024. We do not perform exhaustive hyperparameter tuning; rather, our goal is to showcase the benefits of using RelGT in place of GNNs within the RDL framework. As shown in our ablation of the multi-element tokenization and global module in RelGT (Table 2), and context size (Figure 4), careful tuning may further improve performance across different tasks.
In addition to the HeteroGNN baseline used in RDL, we report results for two variants of the Heterogeneous Graph Transformer (HGT) [20] to highlight the advantages of RelGT over existing GT models. Notably, many GTs, such as GraphGPS [41], are not directly applicable to heterogeneous graphs. Therefore, we adopt HGT and an enhanced version, HGT+PE, which incorporates Laplacian positional encodings (LapPE). These positional encodings are computed on the sampled subgraphs rather than the entire graph. Additional implementation details are included in Appendix A.3.
Table 1: Test set results on the entity regression and classification tasks in RelBench. Best values are in bold. RDL: HeteroGNN baseline [42], HGT: Heterogeneous GT [20], PE: Laplacian Positional Encodings [10]. Relative gains are expressed as percentage improvement over RDL baseline.
(a) MAE for entity regression. Lower is better
Dataset Task RDL HGT HGT +PE RelGT (ours) % Rel. Gain
rel-f1 driver-position 4.022 4.1598 4.2358 3.9170 2.61
rel-avito ad-ctr 0.041 0.0441 0.0494 0.0345 15.85
rel-event user-attendance 0.258 0.2635 0.2562 0.2502 2.79
rel-trial study-adverse 44.473 43.3253 42.4622 43.9923 1.08
site-success 0.400 0.4374 0.4431 0.3263 18.43
rel-amazon user-ltv 14.313 15.3804 15.9296 14.2665 0.32
item-ltv 50.053 56.1384 55.6211 48.9222 2.26
rel-stack post-votes 0.065 0.0679 0.0680 0.0654 -0.62
rel-hm item-sales 0.056 0.0655 0.0641 0.0536 4.29
(b) AUC for entity classification. Higher is better.
Dataset Task RDL HGT HGT +PE RelGT (ours) % Rel. Gain
rel-f1 driver-dnf 0.7262 0.7142 0.7109 0.7587 4.48
driver-top3 0.7554 0.6389 0.8340 0.8352 10.56
rel-avito user-clicks 0.6590 0.6584 0.6387 0.6830 3.64
user-visits 0.6620 0.6426 0.6507 0.6678 0.88
rel-event user-repeat 0.7689 0.6717 0.6590 0.7609 -1.04
user-ignore 0.8162 0.8348 0.8161 0.8157 -0.06
rel-trial study-outcome 0.6860 0.5679 0.5691 0.6861 0.01
rel-amazon user-churn 0.7042 0.6608 0.6589 0.7039 -0.04
item-churn 0.8281 0.7824 0.7840 0.8255 -0.31
rel-stack user-engagement 0.9021 0.8898 0.8852 0.9053 0.35
user-badge 0.8986 0.8652 0.8518 0.8632 -3.94
rel-hm user-churn 0.6988 0.6773 0.6491 0.6927 -0.87
4.2 Results and Discussion
RelGT improves over GNN in RDL. The experimental results in Tables 1(a) and 1(b) demonstrate that RelGT consistently matches or outperforms the standard GNN baseline used in RDL [42] across multiple datasets and tasks. We observe the largest improvements in rel-trial site-success (18.43%), rel-avito ad-ctr (15.85%), and rel-f1 driver-top3 (10.56%), while on rel-stack user-badge, RelGT performs below the RDL baseline by a margin of -3.94%. For all other tasks, RelGT consistently improves or matches the performance of the baseline GNN. We attribute the overall performance improvement to two key factors: (i) the broader structural coverage enabled by RelGT’s attention mechanisms as described in Section 3.2, and (ii) the fine-grained encodings employed in our tokenization scheme, which are further studied as follows and presented in Table 2.
Subgraph GNN PE is critical in RelGT. In Table 2, we highlight the importance of several components in RelGT by conducting ablation studies. We remove one component at a time while preserving all others, and report the relative performance drop compared to the full RelGT model. Our results show that removing the subgraph GNN (PE), which encodes local subgraph structure (Section 3.1), leads to consistent performance degradation across all tasks. This component proves critical for disambiguating parent-child relationships when full-attention is applied, thanks to the random node features initialization [43, 26]. For instance, without the GNN (PE), products belonging to specific transactions (Figure 1) cannot be effectively captured, even when other encodings remain.
Global module can bring gains depending on the task. In the same Table 2, our results of removing the global attention to the learnable centroids (Section 3.2) reveal task-dependent patterns that align with the findings reported in [30, 11]. For some tasks, such as rel-trial site-success, removing the attention to the centroids tokens leads to a substantial performance drop (-19.08%), indicating that the global database-wide context provides crucial information beyond the local neighborhood. However, for certain tasks such as rel-avito user-clicks, removing the global module actually improves performance (7.79% relative gain), suggesting that for some prediction targets, local information is sufficient, and the global context might introduce noise. These mixed results highlight the complementary nature of local and global information in relational graphs, with the latter being optional depending on the task.
Ablation of other encodings. The remaining ablations in Table 2 reveal mixed results across different components. While removing explicit fine-grained encodings (node type, hop distance, and relative time) degrades performance on some tasks, it improves performance on others. For tasks with specific temporal dependencies (as detailed in the Appendix A.1), our current temporal encodings may inadvertently introduce noise. Similarly, for node type and hop distance encodings, their information might already be partially captured by other model components. Despite these variations, the full RelGT model still shows consistently superior results when averaged across all tasks. However, our findings suggest that RelGT’s performance could be further enhanced by careful tuning of these encoding components based on their task-specific importance. In particular, additional improvements can be achieved by incorporating more effective temporal encoding methods [7, 21, 23].
Table 2: Relative drop (%) in performance in RelGT after removing a model component. Negative scores suggest the component is critical in RelGT, and vice-versa. Full results in Table 7.
Dataset Task No Global Module No GNN PE No Node Type No Hop Distance No Relative Time
rel-avito ad-ctr $-6.00$ $-1.14$ $-7.14$ $-3.43$ $-9.14$
rel-avito user-clicks $7.85$ $-15.15$ $5.01$ $5.77$ $8.37$
rel-avito user-visits $-0.35$ $-2.38$ $-0.11$ $0.39$ $-0.75$
rel-event user-ignore $-1.30$ $0.12$ $-0.11$ $0.66$ $-0.09$
rel-trial study-outcome $-2.14$ $-1.72$ $3.74$ $-0.43$ $2.48$
rel-trial site-success $-19.01$ $-9.17$ $-2.88$ $-21.49$ $-0.71$
rel-amazon user-churn $-0.64$ $-0.78$ $0.16$ $0.06$ $-2.20$
rel-hm item-sales $-9.33$ $-17.35$ $-12.69$ $0.93$ $-77.24$
Average $-3.87$ $-5.95$ $-1.75$ $-2.19$ $-9.91$
Figure 4: Left: Epoch runtime comparison of HGT [20] and HGT+PE, with Laplacian PE (see Figure 5 for all tasks). The red portion shows the additional time consumed by the precomputation of Laplacian PE against the base HGT time (blue). Right: Ablation for different $K$ values as the local context size in RelGT. Results using $K=300$ serve as the baseline (100% performance), with $K=100$ and $K=500$ runs measured as % of performance relative to $K=300$.
HGT, a GT baseline, underperforms with significant computational overhead. As shown in Tables 1(a) and 1(b), HGT [20] underperforms compared to the HeteroGNN baseline of RDL [42] across most tasks, with only two exceptions: rel-trial study-adverse and rel-event user-ignore. Notably, the integration of Laplacian eigenvectors as PEs in HGT improves performance in just 5 out of 21 tasks. Moreover, as illustrated in Figure 4, the computational overhead required for precomputing the Laplacian PEs substantially increases per-epoch runtime across various tasks. These empirical findings clearly reveal the difficulties of directly applying existing GT architectures to relational entity graphs, emphasizing the importance and need for our contributions with RelGT.
Local context size K. In our main RelGT experiments, we set the local context size at $300$ nodes (Section 3.1), however, we study its variability in Figure 4 for context sizes $K\in\{100,300,500\}$. Although $K=300$ generally produces the best results, optimal values vary across specific tasks. For instance, rel-avito ad-ctr benefits from a larger context size, whereas rel-trial study-outcome achieves better performance with a smaller context window. These findings suggest that RelGT’s performance could be further enhanced by task-specific tuning of the context size, allowing for better model expressivity based on the structural characteristics of each dataset.
5 Conclusion
In this work, we introduce the first Graph Transformer designed specifically for relational entity graphs: the Relational Graph Transformer RelGT. It addresses key challenges faced by existing models, such as incorporating heterogeneity, temporality, and comprehensive structural modeling within a unified GT framework. RelGT represents nodes as multi-element tokens enriched with fine-grained graph context and combines local attention over sampled subgraph tokens with global attention to learnable centroids, enabling effective representation learning on relational data. Experiments on the RelBench benchmark show that RelGT consistently outperforms GNN and GT baselines across multiple tasks. Moreover, our analysis highlights the critical role of subgraph-based positional encodings as a lightweight and effective alternative to traditional graph positional encodings. This work establishes RelGT as a powerful architecture for relational deep learning and opens new avenues for advancing and scaling such architectures toward foundation models tailored for relational data.
Acknowledgments and Disclosure of Funding
We thank Eric Chen, Shenyang Huang and Fang Wu for their helpful feedbacks and members of the Stanford SNAP group for their suggestions during the project. We also gratefully acknowledge the support of NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, Hitachi, and SAP. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding entities.
Appendix A Appendix
A.1 Benchmark Details
Table 3: Dataset and task statistics from RelBench used for our evaluation.
Dataset Task Task type #Rows of training table #Unique %train/test
Train Validation Test Entities Entity Overlap
rel-amazon user-churn classification 4,732,555 409,792 351,885 1,585,983 88.0
item-churn classification 2,559,264 177,689 166,842 416,352 93.1
user-ltv regression 4,732,555 409,792 351,885 1,585,983 88.0
item-ltv regression 2,707,679 166,978 178,334 427,537 93.5
rel-avito user-clicks classification 59,454 21,183 47,996 66,449 45.3
user-visits classification 86,619 29,979 36,129 63,405 64.6
ad-ctr regression 5,100 1,766 1,816 4,997 59.8
rel-event user-repeat classification 3,842 268 246 1,514 11.5
user-ignore classification 19,239 4,185 4,010 9,799 21.1
user-attendance regression 19,261 2,014 2,006 9,694 14.6
rel-f1 driver-dnf classification 11,411 566 702 821 50.0
driver-top3 classification 1,353 588 726 134 50.0
driver-position regression 7,453 499 760 826 44.6
rel-hm user-churn classification 3,871,410 76,556 74,575 1,002,984 89.7
item-sales regression 5,488,184 105,542 105,542 105,542 100.0
rel-stack user-engagement classification 1,360,850 85,838 88,137 88,137 97.4
user-badge classification 3,386,276 247,398 255,360 255,360 96.9
post-votes regression 2,453,921 156,216 160,903 160,903 97.1
rel-trial study-outcome classification 11,994 960 825 13,779 0.0
study-adverse regression 43,335 3,596 3,098 50,029 0.0
site-success regression 151,407 19,740 22,617 129,542 42.0
In this section, we include the details on the datasets and the tasks in RelBench [42] which we use for our evaluation. RelBench consists of 7 datasets from diverse relational database domains, including e-commerce, clinical records, social networks, and sports, among others. These datasets are curated from their respective source domains and consist a wide range of sizes, from 1.3K to 5.4M records in the training set for the prediction tasks, with a total of 47M training records. For each dataset, multiple predictive tasks are defined, such as predicting a user’s engagement with an advertisement within the next four days or determining whether a clinical trial will achieve its primary outcome within the next year. In total, RelBench has 30 tasks across the 7 datasets, covering entity classification, entity regression, and recommendation. For our evaluation, we focus on 21 tasks on entity classification and regression as RelGT primarily serves as a node representation learning model in RDL. We exclude recommendation tasks in this work since they involve specific considerations, such as identifying target nodes [54] or using pair-wise learning architectures [55] and using RelGT trivially in RDL is sub-optimal. We detail the dataset and task statistics in Table 3.
A.1.1 Datasets
rel-amazon.
The Amazon E-commerce dataset consists of product details, user information, and review interactions from Amazon’s platform, including metadata like pricing and categories, along with review ratings and content.
rel-avito.
Avito’s marketplace dataset contains search queries, advertisement characteristics, and contextual information from this major online trading platform that facilitates transactions across various categories including real estate and vehicles.
rel-event.
The Event Recommendation dataset from Hangtime mobile app tracks users’ social planning, capturing interactions, event details, demographic data, and social connections to reveal how relationships impact user behavior.
rel-f1.
The F1 dataset provides comprehensive Formula 1 racing information since 1950, documenting drivers, constructors, manufacturers, and circuits with detailed records of race results, standings, and specific data on various racing sessions and pit stops.
rel-hm.
H&M’s dataset contains customer-product interactions from their e-commerce platform, featuring customer demographics, product descriptions, and purchase histories.
rel-stack.
The Stack Exchange dataset documents activity from this network of Q&A websites, including user biographies, posts, comments, edits, votes, and question relationships where users earn reputation through contributions.
rel-trial.
The clinical trial dataset from the AACT initiative has study protocols and outcomes, containing trial designs, participant information, intervention details, and results metrics, serving as a key resource for medical research.
A.1.2 Tasks
The following entity classification and regression tasks are defined in RelBench for the above datasets.
  1. 1.
    rel-amazon
    1. (a)
      user-churn: Predict whether a user will discontinue reviewing products within the next three months.
    2. (b)
      item-churn: Predict if a product will have no reviews in the next three months.
    3. (c)
      user-ltv: Estimate the total monetary value of merchandise in dolloar that a user will purchase and review within the next three months.
    4. (d)
      item-ltv: Estimate the total monetary value of purchases and reviews a product will receive during the next three months.
  2. 2.
    rel-avito
    1. (a)
      user-visits: Predict if a user will engage with several (advertisements) ads within the upcoming four days.
    2. (b)
      user-clicks: Predict whether a user will interact with multiple ads through clicking within the upcoming four days.
    3. (c)
      ad-ctr: Estimate the interaction probability for an ad, assuming it receives an interaction within four days.
  3. 3.
    rel-event
    1. (a)
      user-attendance: Estimate the number of of events a user will confirm attendance to (RSVP yes or maybe) within the upcoming seven days.
    2. (b)
      user-repeat: Predict whether a user will join an event (RSVP yes or maybe) within the upcoming seven days, provided they attended in an event during the previous fourteen days.
    3. (c)
      user-ignore: Predict whether a user will disregard or ignore more than two events invitations within the upcoming seven days.
  4. 4.
    rel-f1
    1. (a)
      driver-dnf: Predict if a driver will not finish a race within the upcoming month.
    2. (b)
      driver-top3: Determine if a driver will achieve a top-three qualifying position in a race within the upcoming month.
    3. (c)
      driver-position: Estimate a driver’s average finishing placement across all races in the upcoming two months.
  5. 5.
    rel-hm
    1. (a)
      user-churn: Predict whether a customer will not perform any transactions in the upcoming week.
    2. (b)
      item-sales: Estimate total revenue generated by a product in the upcoming week.
  6. 6.
    rel-stack
    1. (a)
      user-engagement: Predict whether a user will contribute through voting, posting, or commenting within the upcoming three months.
    2. (b)
      user-badge: Predict whether a user will secure a new badge within the upcoming three months.
    3. (c)
      post-votes: Estimate the number of votes a user’s post will accumulate over the upcoming three months.
  7. 7.
    rel-trial
    1. (a)
      study-outcome: Predict whether a clinical trial will achieve its principal outcome within the upcoming year.
    2. (b)
      study-adverse: Estimate the number of patients who will experience significant adverse effects or mortality in a clinical trial over the upcoming year.
    3. (c)
      site-success: Estimate the success rate of a clinical trial site in the upcoming year.
A.2 Node initialization for Subgraph GNN PE in RelGT
As described in Section 3.1, we employ a lightweight GNN PE to capture local graph structures that cannot be represented by other elements of the token, particularly the parent-child relationships among nodes in the local subgraph. The GNN is implemented as:
$h_{\text{pe}}(v_{j})=\text{GNN}(A_{\text{local}},Z_{\text{random}})_{j}\in\mathbb{R}^{d}$ (10)
where $\text{GNN}(\cdot,\cdot)_{j}$ is a lightweight GNN applied to the local subgraph, yielding the encoding for node $v_{j}$. Here, $A_{\text{local}}\in\mathbb{R}^{K\times K}$ represents the adjacency matrix of the sampled subgraph containing $K$ nodes, and $Z_{\text{random}}\in\mathbb{R}^{K\times d_{\text{init}}}$ denotes randomly initialized node features for the GNN (with $d_{\text{init}}$ as the initial feature dimension). In RelGT, we set $d_{\text{init}}=1$.
The randomly initialized node features ($Z_{\text{random}}$) provide enhanced properties as discussed in Section 3.1. We investigate the alternative approach of using Laplacian PE ($Z_{\text{LapPE}}$) computed over the subgraph instead of random initialization and report these results in Table 4. For these results, we utilized a positional encoding dimension size of $4$. Our findings indicate that $Z_{\text{LapPE}}$ consistently underperforms compared to $Z_{\text{random}}$, while also introducing additional computational overhead ranging from 1.02$\times$ to 3.38$\times$ across the 8 selected tasks in our study. This shows the challenges of using existing PEs such as Laplacian PE in relational entity graphs and signify the use of GNN PE as part of RelGT’s tokenization strategy.
Table 4: Study of node initialization in Subgraph GNN PE. Relative drop is expressed as percentage drop of using $Z_{\text{LapPE}}$ vs. $Z_{\text{random}}$ and runtime ratio compares the time for $Z_{\text{LapPE}}$ vs. $Z_{\text{random}}$.
Performance Epoch time (m)
Dataset Task (# train) MAE $\downarrow$ $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Rel Drop $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Runtime Ratio
rel-avito ad-ctr Test 0.035 0.0369 -5.43 0.76 2.57 3.38
Val 0.0314 0.0314
rel-trial site-success Test 0.326 0.3452 -5.89 32.88 36.09 1.1
Val 0.359 0.3683
rel-hm item-sales Test 0.0536 0.0573 -6.9 49.26 53.8 1.09
Val 0.0627 0.0667
Dataset Task (# train) AUC $\uparrow$ $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Rel Drop $Z_{\text{random}}$ $Z_{\text{LapPE}}$ Runtime Ratio
rel-avito user-clicks Test 0.607 0.583 -3.95 6.42 7.43 1.16
Val 0.656 0.6564
user-visits Test 0.664 0.6626 -0.21 9.26 10.50 1.13
Val 0.699 0.7002
rel-event user-ignore Test 0.8 0.7988 -0.15 1.85 2.77 1.5
Val 0.881 0.8916
rel-trial study-outcome Test 0.674 0.6532 -3.09 1.41 1.52 1.08
Val 0.689 0.6719
rel-amazon user-churn Test 0.7039 0.7044 0.07 168.00 170.55 1.02
Val 0.7036 0.7036
A.3 HGT Baseline
In the main experiments (Section 4), we use the Heterogeneous Graph Transformer (HGT) [20] as a graph transformer (GT) baseline, and report results for two variants to demonstrate the advantages of RelGT over existing GT models. Specifically, we consider the standard HGT model and an enhanced version, HGT+PE, which incorporates Laplacian positional encodings (LapPE). These positional encodings are computed on sampled subgraphs rather than the full graph.
For implementation, we use the HGTConv layer from PyTorch Geometric [14] and integrate it into the RDL pipeline [42] by replacing the default GNN module. Both variants use 4 attention heads and 2 layers, similar to the configuration of the GNN module in RDL, with residual connections and layer normalization applied between layers. For the HGT+PE variant, we use LapPE of dimension 4 for all tasks, except for rel-amazon item-ltv and rel-hm item-sales, where we use dimension 2. Notably, because the relational entity graphs are heterogeneous, the Laplacian positional encodings is computed multiple times for each node type, unlike the original homogeneous setting for which LapPE was designed [10].
In addition to the main results in Table 1(b), we report per-epoch runtimes in Figure 5 and Table 5. We observe a significant computational overhead from precomputing Laplacian positional encodings, with slowdowns ranging from 1.8× to 8.62×, highlighting the challenge of directly applying existing graph PE techniques as is to relational entity graphs, and signifying the contributions of RelGT.
Figure 5: Runtime Comparison of HGT and HGT+PE baseline. Adding the Laplacian Positional Encoding increases computational overhead, with penalties on average training time per epoch. The overhead for PE reaches up to 761% relative to the training time of HGT on the same dataset.
Table 5: Relative performance drop (%) when position encoding (PE) is removed from HGT+PE models and average training time per epoch of HGT and HGT+PE. Negative scores suggest the PE is critical, and vice-versa. HGT+PE consistently requires more training time per epoch compared to HGT without PE across all datasets.
Dataset Task No PE HGT(s) HGT+PE(s)
rel-f1 driver-position $1.79$ 1.47 4.56
rel-avito ad-ctr $10.73$ 1.63 4.00
rel-event user-attendance $-2.85$ 4.36 37.57
rel-trial study-adverse $-2.03$ 9.72 15.02
rel-trial site-success $1.29$ 45.73 63.41
rel-amazon user-ltv $3.45$ 73.59 106.21
rel-amazon item-ltv $-0.93$ 73.68 110.33
rel-stack post-votes $0.15$ 191.23 528.25
rel-hm item-sales $-2.18$ 94.66 135.05
rel-f1 driver-dnf $0.46$ 2.54 6.84
rel-f1 driver-top3 $-23.39$ 0.38 1.38
rel-avito user-clicks $3.08$ 11.09 24.66
rel-avito user-visits $-1.24$ 17.16 40.07
rel-event user-repeat $1.93$ 1.35 5.16
rel-event user-ignore $2.29$ 4.49 651.10
rel-trial study-outcome $-0.21$ 4.09 4.83
rel-amazon user-churn $0.29$ 78.56 115.53
rel-amazon item-churn $-0.20$ 75.51 152.06
rel-stack user-engagement $0.52$ 175.16 356.07
rel-stack user-badge $1.57$ 153.68 212.21
rel-hm user-churn $4.34$ 77.73 127.04
Average $-0.05$ 52.28 128.64
A.4 Detailed Results
In Table 6, we report the full results of different configurations we tuned for RelGT, particularly on the smaller datasets with lesser than a million training nodes. Table 7 provides the full scores for the RelGT component study in Table 2, while Table 8 provides the supporting results for Figure 4. Finally, we provide the elaborated version of the Tables 1(a) and 1(b) in Tables 9 and 9, respectively.
A.5 Resource Information.
We implement RelGT using PyTorch framework [39], PyTorch Geometric framework [14] and adapt the codebase of relational deep learning [42] https://github.com/snap-stanford/relbench. All our experiments are conducted on an NVIDIA A100 GPU server with 8 GPU nodes.
Table 6: RelGT results using $L\in{1,4,8}$ and dropout $\in{0.3,0.4,0.5}$ for the smaller datasets with less than a million training nodes.
Dataset Task (# train) MAE $\downarrow$ L1 0.3 L1 0.4 L1 0.5 L4 0.3 L4 0.4 L4 0.5 L8 0.3 L8 0.4 L8 0.5
rel-f1 driver-position (7k) Test 4.942 5.6431 3.917 4.6316 4.0851 4.0042 5.5273 5.5569 4.6085
Val 3.1897 3.1817 3.3257 3.1046 3.3352 3.1276 3.1589 3.2907 3.1843
rel-avito ad-ctr (5k) Test 0.0358 0.0352 0.0345 0.035 0.0366 0.038 0.0354 0.0358 0.0356
Val 0.0322 0.0313 0.0314 0.0314 0.0322 0.0335 0.0317 0.0322 0.0324
rel-event user-attendance (19k) Test 0.2635 0.2595 0.2635 0.2502 0.2543 0.2584 0.2635 0.2637 0.2635
Val 0.2618 0.2558 0.2618 0.2548 0.2534 0.253 0.2618 0.2599 0.2618
rel-trial study-adverse (43k) Test 44.8553 44.2260 44.848 44.8893 44.4310 43.9923 44.2245 44.5878 44.5013
Val 46.3538 46.3193 46.2056 46.1031 45.9498 46.2148 46.1804 46.1381 46.4332
site-success (151k) Test 0.3490 0.3652 0.3830 0.4019 0.386 0.3262 0.3783 0.3431 0.3644
Val 0.3493 0.3455 0.3550 0.3771 0.392 0.3593 0.3848 0.3643 0.3669
Dataset Task (# train) AUC $\uparrow$ L1 0.3 L1 0.4 L1 0.5 L4 0.3 L4 0.4 L4 0.5 L8 0.3 L8 0.4 L8 0.5
rel-f1 driver-dnf (11k) Test 0.7434 0.7587 0.7521 0.7587 0.745 0.6957 0.7349 0.7393 0.741
Val 0.6877 0.6761 0.6896 0.6804 0.6762 0.6768 0.6702 0.6803 0.6865
driver-top3 (1k) Test 0.7845 0.8203 0.8 0.8171 0.8157 0.8352 0.7871 0.8217 0.8222
Val 0.7775 0.783 0.7764 0.7841 0.79 0.7958 0.7893 0.7847 0.7829
rel-avito user-clicks (59k) Test 0.6524 0.6233 0.6212 0.6067 0.5893 0.596 0.6245 0.683 0.6507
Val 0.6649 0.6616 0.6501 0.6564 0.6608 0.6579 0.6587 0.6649 0.6648
user-visits (86k) Test 0.6627 0.6663 0.665 0.6615 0.6584 0.6642 0.6647 0.6678 0.664
Val 0.7005 0.6993 0.7001 0.6954 0.6958 0.699 0.6995 0.7024 0.7011
rel-event user-repeat (3k) Test 0.6981 0.7403 0.7452 0.7563 0.7236 0.7432 0.7609 0.7316 0.7418
Val 0.7172 0.7386 0.7319 0.7245 0.7207 0.736 0.7285 0.7209 0.7064
user-ignore (19k) Test 0.8006 0.802 0.7986 0.799 0.787 0.8002 0.7956 0.8076 0.8157
Val 0.8739 0.8721 0.8729 0.878 0.8731 0.881 0.8757 0.8801 0.8868
rel-trial study-outcome (11k) Test 0.6808 0.6753 0.6837 0.6488 0.6818 0.6744 0.6861 0.6562 0.6649
Val 0.6815 0.6792 0.6751 0.6737 0.676 0.689 0.6678 0.6746 0.6768
Table 7: Relative drop (%) in performance in RelGT after removing a model component. Negative scores suggest the component is critical in RelGT, and vice-versa.
Dataset Task (# train) MAE $\downarrow$ RelGT (Full) RelGT (No Global) % Rel. Drop RelGT (No GNN) % Rel. Drop RelGT (No Type) % Rel. Drop RelGT (No Hop) % Rel. Drop RelGT (No Time) % Rel. Drop
rel-avito ad-ctr Test 0.0350 0.0371 -6.0 0.0354 -1.14 0.0375 -7.14 0.0362 -3.43 0.0382 -9.14
Val 0.0314 0.0323 0.0315 0.0328 0.0322 0.0337
rel-trial site-success Test 0.3262 0.3882 -19.01 0.3561 -9.17 0.3356 -2.88 0.3963 -21.49 0.3285 -0.71
Val 0.3593 0.3342 0.3637 0.3655 0.3614 0.3615
rel-hm item-sales Test 0.0536 0.0586 -9.33 0.0629 -17.35 0.0604 -12.69 0.0531 0.93 0.095 -77.24
Val 0.0627 0.0676 0.073 0.0696 0.0623 0.1025
Dataset Task (# train) AUC $\uparrow$ RelGT (Full) RelGT (No Global) % Rel. Drop RelGT (No GNN) % Rel. Drop RelGT (No Type) % Rel. Drop RelGT (No Hop) % Rel. Drop RelGT (No Time) % Rel. Drop
rel-avito user-clicks Test 0.6067 0.6543 7.85 0.5148 -15.15 0.6371 5.01 0.6417 5.77 0.6575 8.37
Val 0.6564 0.6496 0.6551 0.6559 0.6482 0.6579
user-visits Test 0.6642 0.6619 -0.35 0.6484 -2.38 0.6635 -0.11 0.6668 0.39 0.6592 -0.75
Val 0.699 0.6892 0.6879 0.6991 0.7016 0.7005
rel-event user-ignore Test 0.8002 0.7898 -1.3 0.8012 0.12 0.7993 -0.11 0.8055 0.66 0.7995 -0.09
Val 0.881 0.8575 0.8637 0.8873 0.8852 0.8789
rel-trial study-outcome Test 0.6744 0.66 -2.14 0.6628 -1.72 0.6996 3.74 0.6715 -0.43 0.6911 2.48
Val 0.689 0.664 0.6775 0.6728 0.6705 0.6578
rel-amazon user-churn Test 0.7039 0.6994 -0.64 0.6984 -0.78 0.705 0.16 0.7043 0.06 0.6884 -2.2
Val 0.7036 0.6994 0.6994 0.7042 0.704 0.6882
Table 8: Ablation of context size $K$ in RelGT.
Dataset Task (# train) MAE $\downarrow$ RelGT K=100 RelGT K=300 RelGT K=500
rel-avito ad-ctr Test 0.0375 0.0374 0.0351
Val 0.0329 0.0319 0.031
rel-trial site-success Test 0.3739 0.3674 0.3842
Val 0.3708 0.372 0.376
rel-hm item-sales Test 0.055 0.0532 0.052
Val 0.0643 0.0619 0.061
Dataset Task (# train) AUC $\uparrow$ RelGT K=100 RelGT K=300 RelGT K=500
rel-avito user-clicks Test 0.6628 0.6491 0.6334
Val 0.6437 0.6622 0.6632
user-visits Test 0.6664 0.6653 0.6627
Val 0.7013 0.701 0.7005
rel-event user-ignore Test 0.7674 0.8105 0.8068
Val 0.8682 0.8853 0.8843
rel-trial study-outcome Test 0.7078 0.6526 0.666
Val 0.6575 0.663 0.6877
rel-amazon user-churn Test 0.7038 0.7054 0.7043
Val 0.7033 0.7044 0.7042
Table 9: Results on the entity regression tasks in RelBench. Lower is better. Best values are in bold. Relative gains are expressed as percentage improvement over RDL baseline.
Dataset Task MAE $\downarrow$ RDL Baseline HGT HGT +PE RelGT (ours) % Rel. Gain
rel-f1 driver-position Test 4.022 4.1598 4.2358 3.9170 2.61
Val 3.193 3.3517 2.9894 3.3257
rel-avito ad-ctr Test 0.041 0.0441 0.0494 0.0345 15.85
Val 0.037 0.0409 0.0456 0.0314
rel-event user-attendance Test 0.258 0.2635 0.2562 0.2502 2.79
Val 0.255 0.2617 0.2574 0.2548
rel-trial study-adverse Test 44.473 43.3253 42.4622 43.9923 1.08
Val 46.290 45.9957 45.7966 46.2148
site-success Test 0.400 0.4374 0.4431 0.3263 18.43
Val 0.401 0.4198 0.4245 0.3593
rel-amazon user-ltv Test 14.313 15.3804 15.9296 14.2665 0.32
Val 12.132 13.1017 13.5599 12.1151
item-ltv Test 50.053 56.1384 55.6211 48.9222 2.26
Val 45.1401 51.2139 50.3468 43.8161
rel-stack post-votes Test 0.065 0.0679 0.0680 0.0654 -0.62
Val 0.059 0.0617 0.0618 0.0592
rel-hm item-sales Test 0.056 0.0655 0.0641 0.0536 4.29
Val 0.065 0.0749 0.0735 0.0627
Table 10: Results on the entity classification tasks in RelBench. Higher is better. Best values are in bold. Relative gains are expressed as percentage improvement over RDL baseline.
Dataset Task AUC $\uparrow$ RDL Baseline HGT HGT +PE RelGT (ours) % Rel. Gain
rel-f1 driver-dnf Test 0.7262 0.7142 0.7109 0.7587 4.48
Val 0.7136 0.7678 0.7318 0.6804
driver-top3 Test 0.7554 0.6389 0.8340 0.8352 10.56
Val 0.7764 0.6659 0.6079 0.7958
rel-avito user-clicks Test 0.6590 0.6584 0.6387 0.6830 3.64
Val 0.6473 0.5977 0.5656 0.6649
user-visits Test 0.6620 0.6426 0.6507 0.6678 0.88
Val 0.6965 0.6696 0.6732 0.7024
rel-event user-repeat Test 0.7689 0.6717 0.6590 0.7609 -1.04
Val 0.7125 0.6247 0.5974 0.7285
user-ignore Test 0.8162 0.8348 0.8161 0.8157 -0.06
Val 0.9170 0.8896 0.8940 0.8868
rel-trial study-outcome Test 0.6860 0.5679 0.5691 0.6861 0.01
Val 0.6818 0.5985 0.5925 0.6678
rel-amazon user-churn Test 0.7042 0.6608 0.6589 0.7039 -0.04
Val 0.7045 0.6639 0.6622 0.7036
item-churn Test 0.8281 0.7824 0.7840 0.8255 -0.31
Val 0.8239 0.7845 0.7846 0.8220
rel-stack user-engagement Test 0.9021 0.8898 0.8852 0.9053 0.35
Val 0.9059 0.8914 0.8847 0.9033
user-badge Test 0.8986 0.8652 0.8518 0.8632 -3.94
Val 0.8886 0.8760 0.8691 0.8741
rel-hm user-churn Test 0.6988 0.6773 0.6491 0.6927 -0.87
Val 0.7042 0.6814 0.6502 0.6988